# Close-Fitting Dressing Assistance Based on State Estimation of Feet and Garments with Semantic-based Visual Attention

### v1

@@ Abstract
As the population continues to age, a shortage of caregivers is expected in the future. Dressing assistance, in particular, is crucial for opportunities for social participation. Especially dressing close-fitting garments, such as socks, remains challenging due to the need for fine force adjustments to handle the friction or snagging against the skin, while considering the shape and position of the garment. This study introduces a method uses multi-modal information including not only robot’s camera images, joint angles, joint torques, but also tactile forces for proper force interaction that can adapt to individual differences in humans. Furthermore, by introducing semantic information based on object concepts, rather than relying solely on RGB data, it can be generalized to unseen feet and background. In addition, incorporating depth data helps infer relative spatial relationship between the sock and the foot. To validate its capability for semantic object conceptualization and to ensure safety, training data were collected using a mannequin, and subsequent experiments were conducted with human subjects. In experiments, the robot successfully adapted to previously unseen human feet and was able to put socks on 10 participants, achieving a higher success rate than Action Chunking with Transformer and Diffusion Policy. These results demonstrate that the proposed model can estimate the state of both the garment and the foot, enabling precise dressing assistance for close-fitting garments.
@@ I INTRODUCTION
In garment manipulation, representing the state of the garment and the human body, as well as modeling their dynamics, is challenging due to occlusions and garment deformation. In addition, the safety of the humans should be ensured. Recent studies have tracked these challenges based on multimodal information using deep learning techniques [1, 2, 3, 4]. Previous research on dressing assistance has focused on garments, such as shirts and trousers, that have wide openings and allow for considerable misalignment between the garment and the body during movement [5]. It is still difficult for a robot to make fine adjustments to the force applied when dressing, while accounting for the garment’s shape and its friction/hooking with the skin. Close-fitting dressing assistance such as putting on socks has not been fully realized. Furthermore, individual physique variations are also one of the factors that complicate dressing assistance, highlighting the need for the generalizability of robotic manipulation.
@@
  • Establishing a imitation learning method for policy of close-fitting garment manipulation with proper force interaction adapting to individual differences in humans
  • Evaluating generalizability of the proposed method, which was trained on a mannequin’s foot for safety and applied to an untrained real human feet.
@@ III Problem formulation and assumptions
Close-fitting garment, such as socks, remains a significant challenge due to their complex deformation and close contact with human skin. During dressing assistance, the garment often gets caught on the body, making it challenging to dress the human. Accurate force interaction, which guides the garment along the body with appropriate force while avoiding the application of excessive force to the human with physical limitations, is fundamental. Multimodal integration empowered by deep-learning technologies is a key factor to address complex force interaction, considering fabric elasticity, friction at the sock-foot interface, and the directional control of the pulling force. However, when it comes to the use of imitation learning, the human difference inhibits the accurate force interaction to the human. On the other hand, using reinforcement learning has the challenge of simulating accurate interactive force. In this study, we formulate the problem of planning and executing the sophisticated action of changing socks by a humanoid robot in a generalized manner that can adapt to individual differences in physique, such as human foot size, shape, color, and flexibility.
According to the two visual streams hypothesis, the human visual system processes information through two distinct pathways: the dorsal stream for spatial awareness and action guidance, and the ventral stream for object recognition and semantic understanding [23, 24]. Recent findings emphasize the importance of interactions between these two pathways, especially in the context of complex object manipulation such as skilled grasping. In this study, we hypothesize that enabling a robot to assist in close-fitting garment-wearing tasks requires a combined understanding of both the object’s semantic features and its spatial characteristics (e.g., depth, orientation relative to the human foot). Based on the ventral-dorsal interaction model, we posit that integrating semantic segmentation with depth estimation allows the system to plan and execute more adaptive and precise manipulations. To model the ventral and dorsal processing streams computationally, we utilize the Segment Anything 2 (SAM2) model [25] for semantic segmentation, which corresponds to the ventral stream’s role in object identification. For estimating depth, we adopt the Depth Anything Model (DAM) [26], which mimics dorsal stream processing for spatial localization and motor planning. We expect that this dual-stream inspired integration will improve the robot’s ability to handle deformable, non-rigid objects in interaction with the human body, especially under conditions requiring fine motor adjustments and adaptive contact.
@@ IV-A Semantic extraction of garment and human foot
SAM2 and DAM are applied to estimate the state of socks and the human foot. SAM2 generates semantic mask images of socks and feet, and DAM generates depth images from a monocular RGB image. In general, segmentation models are vulnerable to external factors such as object deformation over time, occlusion, and varying lighting conditions[25]. During the task of sock dressing, various types of occlusions frequently occur, such as the sock itself covering the foot or the robot’s arms obscuring the target area. Furthermore, we utilize the SAM2 to mitigate the influence of variations in the color and appearance of the target object (e.g., socks or feet), as well as differences in the background surrounding the object.
The images generated by SAM2 are black and white, with brightness values of 0 and 255, where the masked regions have a brightness value of 255. In parallel, the depth images of the images acquired from the robot’s camera are obtained using DAM[26]. The robot lacks perception in the depth direction when relying solely on two-dimensional images, so we employ DAM to infer the relative spatial relationship between the sock and the toes during insertion. Additionally, DAM facilitates estimating the vertical displacement of the arm required to successfully guide the sock over the heel region. Then, the mask depth images of the socks and feet are obtained by embedding the depth values of the depth image regions corresponding to the masked regions with a brightness value of 255. By adding depth information (z value) of the pixel value of the key points on the 2D coordinates of the image (x coordinate, y coordinate), 3D information on the image (x coordinate, y coordinate, z value) is output.
@@
  • $D\in\mathbb{R}^{H\times W}$: The depth map obtained from the Depth Anything model.
@@ IV-B Visual and somatosensory attention mechanism
As a part of visual attention mechanism, Spatial Attention Recurrent Neural Network (SARNN) is applied for robustness to background change [29]. The image encoder of SARNN compresses and extracts features from visual information, and the attention encoder extracts spatial attention points (coordinates) using the mask images as input. Multiple attention encoders enable the simultaneous estimation of attention points for multiple objects. To further enhance spatial understanding, we incorporate attention that considers depth. For each feature channel, a spatial softmax is applied to generate an attention map, which is used to weight the 2D spatial coordinates and depth map to calculate the expected 3D keypoints. This allows us to focus attention not only on 2D salient regions but also on the 3D geometric context, enabling robust keypoint localization even in occlusion and unknown situations.
@@ IV-C Hierarchical LSTM
This feedback vector is split into modality-specific components:
@@ IV-D Training and inference phase
where $L_{angle}$ represents the mean squared error (MSE) of the joint angles; $L_{torque}$ is the MSE of the joint torques; $L_{tactile}$ is the MSE of the tactile information of the thumbs; $L_{img}$ is the MSE of the predicted image; $L_{pt}$ is the MSE of the attention point $pt=(x_{t},y_{t})$; $\alpha$, $\beta$, $\gamma$, $\delta$, $\epsilon$ represent the loss contributions, and $\alpha$ and $\epsilon$ were set to 0.1. $\beta$ was set to 1.5. $\gamma$ was set to 1.0. $\delta$ was set to 0.2. Adam was used as the optimization model, and the model was trained for a total of 10,000 epochs.
@@ V-A Robot constitutions
In this study, we used a humanoid robot (Fig. 3) developed by Tokyo Robotics Co., Ltd. based on Torobo [30]. It integrates various sensing functions such as torque sensors, tactile sensors, and visual sensors, making it a platform for studying robot interaction with the real world. This makes the robot adept at manipulating various objects in daily activities. In addition, the control system is highly versatile, supporting various modes such as position, force, and impedance control, and parameters such as mass, damping, and spring can be customized. This adaptability allows the robot to perform tasks while improving safety enhancing safety during contact with humans. Each arm of the robot has seven joints, and impedance controllers are used for the joints in the hand that may come into contact with the robot to soften them and improve safety. In addition, in this study, the robot needs to reach low positions at its feet, but the robot is equipped with hardware that allows it to reach down toward lower positions by bending its waist.
@@ V-B Demonstration with PS controller
The robot was teleoperated using a PS controller. Hand positions were updated by adding xyz-direction displacements generated from the controller inputs, and the target joint angles were computed via inverse kinematics. Two movement patterns were used: (1) both arms moving in parallel with identical displacements, and (2) arms moving alternately in opposite directions. The parallel pattern, being intuitive and easy to reproduce, was primarily used for generating consistent demonstration data, while the alternate mode was introduced to handle more diverse motion patterns caused by friction or snagging.
@@ V-C Environment and dataset
In this study, a kit mannequin made in Japan was used as a training subject. The specifications of the mannequin are 170 cm in height, 40 cm in shoulder width, and 19 cm in foot size. The exterior is made of urethane and the interior is made of polystyrene foam. The training environment is shown in Fig. 5. The mannequin was seated in a chair and foot was on a footrest. The seat height of the chair and the footrest was set to 45 cm. We assume a scenario in which a robot puts on socks for a person whose legs are out of the bed and whose feet are off the floor. The mannequin’s body is light, thus the leg joints bend easily. The mannequin’s feet are fixed with strings during training to apply the movements to real humans. Putting on socks requires balance ability, back and lower limb strength, and flexibility, which makes it a difficult movement for people with physical limitations. The requirement for humans to participate in advance for dataset collection places a burden on them. Therefore, a mannequin was used in the training phase for safety, and the adaptability of the model to humans was evaluated.
Furthermore, as shown in Fig. 4, since the way people elevate their feet varies from person to person, we set the mannequin’s foot angle at three different positions—30°, 40°, and 50° from the horizontal plane—to account for these individual differences. For each angle, we collected and used 4 samples, resulting in a total of 12 data samples for training. The algorithm was set to a frequency of 20 Hz, resulting in 210 time steps per dataset. The dataset includes the robot’s vision, the joint angles and torques of both arms’ joints 1-7, and the tactile information of the thumb (FSR), as shown in Fig. 3. The initial joint angles of the head were set to 0°, 37°, and 0° for head/joint1, head/joint2, and head/joint3, respectively. Similarly, the initial angles for the torso were set to 0°, -45°, and 100° for torso/joint1, torso/joint2, and torso/joint3, respectively. An impedance controller is applied to joint 7 to ensure safety in scenarios where contact with the human body may occur. The other joints employ position control, as actuation force is required to manipulate the sock effectively during the dressing process. In this study, the tactile information of the thumb is used to grasp a sock with the thumb and recognize the tightness of the sock.
@@ V-D2 Evaluation of generalization capability and robustness
To evaluate the generalization performance and robustness of the proposed model, we conducted experiments on 10 real humans in a real environment and compared it with two existing models. The participants consisted of 10 people with foot sizes ranging from 23.0 cm to 26.5 cm, and we verified the model’s adaptability to environmental changes and individual differences in feet. The experiment was carried out on the basis of approval from the Ethics Review Committee on Human Subject Research.
@@ VI-A Ablation study on model components
As shown in Table I, the proposed model achieved a 100% success rate. Also,removal of DAM or SKNet resulted in moderate performance degradation, while removal of Hierarchical LSTM or SAM2 and DAM caused a dramatic drop in success rate from 100% to 5% and from 100% to 0% (shown in Fig. 6), respectively. These results confirm the critical role of temporal modeling and the importance of spatial and depth-aware attention mechanisms. Without DAM, the model achieved a success rate of 85%. However, one failure case was observed: after the toe was inserted into the sock, when the arm was lowered along the surface of the feet, the toe got caught in the sock, making it difficult to continue movement. We assume that this was because the positional relationship between the sock and the feet in the depth direction was not properly recognized.
@@
Model Variant Success/Trials
Ours 20/20
No DAM 17/20
No SKNet 16/20
No Hierarchical LSTM 1/20
No SAM and DAM 0/20
@@ VI-B Results of generalization capability and robustness
We evaluated the generalization performance of the proposed method in the 10 human participants with foot sizes ranging from 23.0 cm to 26.5 cm under two different visual conditions: trained (seen background) and untrained (unseen background). Performance was compared with two baselines: ACT and DP. Each method was tested in 50 trials per condition.
@@
As summarized in Table III, our method achieved a success rate of 84% (42/50) in the trained background and 74% (37/50) in the untrained background. In contrast, ACT achieved 66% (33/50) and 0% (0/50), respectively, and DP failed to complete the dressing task under both conditions.
Fig. 7 shows the time series of thumb tactile values for different foot sizes. In the proposed method, the tactile values remained stable across foot sizes. In contrast, in the ACT, a sharp increase around 125 time steps reflects the thumb catching on the heel (Fig. 6). These results demonstrate that the proposed method maintains consistent contact manipulation performance adapting to foot size variations, whereas the ACT is more sensitive to size variations.
A scene of attention points moving over the object area is shown in Fig. 8, with time-series semantic mask images. The attention points consistently track the object area. These results demonstrate that the proposed model is robust to unseen environments and individual anatomical variations. Representative test scenes under both known and unknown backgrounds are shown in Fig. 9.
We analyzed the generalization performance of the proposed model and ACT. The box plots for two models are shown in Fig. 10. This box plot shows the distribution of the number of successes out of 5 for 10 participants. We conducted a Wilcoxon signed rank test (nonparametric was confirmed by Shapiro-Wilk test) on the combined number of successes in the seen and unseen backgrounds. The p-value was 0.0008, which is less than 0.01, indicating a statistically significant difference between the proposed model and ACT.
@@
TABLE III: Success rates corresponding to foot sizes.
@@
Figure 8: A scene of attention points moving over the object area. Blue points and red points indicate current and future image attention key points, respectively.
@@
Stability and reproducibility are particularly important in the task of putting on socks. In DP, there are times when the consistency of the behavior is lost due to the probabilistic nature of the behavior generation, and it becomes clear that the movement of reaching the sock to the toes is difficult.
ACT tended to vary the arm trajectory. In an unknown background, ACT collided with the foot even in the first phase of reaching the foot. We assume that this is because the self-attention mechanism in Transformer has low sensitivity to local changes and tends to be unable to fully respond to differences due to the size of the subjects’ feet.
During the user study, there were instances where the subject’s sock became caught on a nail, making it difficult to continue the dressing motion. This necessitated motion replanning, highlighting the inherent complexity and unpredictability involved in assistive sock dressing tasks.
@@
In this paper, we propose a novel method for multimodal imitation learning with a hierarchical LSTM of a humanoid robot assisting close-fitting garment dressing, adapting to individual differences in feet and changes in the operating environment. Visual information is embedded with depth information in a semantic mask image, and the system learns depth information while recognizing the changing shapes of the foot and socks. The model trained on a mannequin’s foot demonstrated higher performance than the baseline (ACT and DP) in a subject experiment with 10 participants, taking into account various foot sizes, foot widths, and flexibility.

### v2

@@ Abstract
As the population continues to age, a shortage of caregivers is expected in the future. Dressing assistance, in particular, is crucial for opportunities for social participation. Especially dressing close-fitting garments, such as socks, remains challenging due to the need for fine force adjustments to handle the friction or snagging against the skin, while considering the shape and position of the garment. This study introduces a method that uses multi-modal information including not only the robot’s camera images, joint angles, joint torques, but also tactile forces for proper force interaction that can adapt to individual differences in humans. Furthermore, by introducing semantic information based on object concepts, rather than relying solely on RGB data, it can be generalized to unseen feet and background. In addition, incorporating depth data helps infer relative spatial relationship between the sock and the foot. To validate its capability for semantic object conceptualization and to ensure safety, training data were collected using a mannequin, and subsequent experiments were conducted with human subjects. In experiments, the robot successfully adapted to previously unseen human feet and was able to put socks on 10 participants, achieving a higher success rate than Action Chunking with Transformer and Diffusion Policy. These results demonstrate that the proposed model can estimate the state of both the garment and the foot, enabling precise dressing assistance for close-fitting garments.
@@ I INTRODUCTION
In garment manipulation, representing the state of the garment and the human body, as well as modeling their dynamics, is challenging due to occlusions and garment deformation. Moreover, the safety of humans should be considered. Recent studies have addressed these challenges by leveraging multimodal information with deep learning techniques [11, 10, 18, 23]. Previous research on dressing assistance has focused on garments, such as shirts and trousers, that have wide openings and allow for considerable misalignment between the garment and the body [32]. It is still difficult for a robot to make fine adjustments to the force applied when dressing, while accounting for the garment’s shape and its friction/hooking with the skin. Close-fitting dressing assistance, such as putting on socks, has not been fully realized. Furthermore, individual physique variations are also one of the factors that complicate dressing assistance, highlighting the need for the generalizability of robotic manipulation.
@@
  • Establishing an imitation learning method for policy of close-fitting garment manipulation with proper force interaction adapting to individual differences in humans
  • Evaluating generalizability of the proposed method, which was trained on a mannequin’s foot for safety and applied to untrained real human feet.
@@ III Problem formulation and assumptions
Close-fitting garments, such as socks, remain a significant challenge due to their complex deformation and close contact with human skin. During dressing assistance, the garment often gets caught on the body, making it challenging to dress the human. Accurate force interaction, which guides the garment along the body with appropriate force while avoiding the application of excessive force to the human with physical limitations, is fundamental. Multimodal integration empowered by deep-learning technologies is a key factor to address complex force interaction, considering fabric elasticity, friction at the sock-foot interface, and the directional control of the pulling force. However, when it comes to the use of imitation learning, individual differences among humans inhibit accurate force interaction. On the other hand, using reinforcement learning has the challenge of simulating accurate interactive force. In this study, we formulate the problem of planning and executing the sophisticated action of putting on socks by a humanoid robot in a generalized manner that can adapt to individual differences in physique, such as human foot size, shape, color, and flexibility.
According to the two visual streams hypothesis, the human visual system processes information through two distinct pathways: the dorsal stream for spatial awareness and action guidance, and the ventral stream for object recognition and semantic understanding [21, 20]. Recent findings emphasize the importance of interactions between these two pathways, especially in the context of complex object manipulation such as skilled grasping. In this study, we hypothesize that enabling a robot to assist in close-fitting garment-wearing tasks requires a combined understanding of both the object’s semantic features and its spatial characteristics (e.g., depth, orientation relative to the human foot). Based on the ventral-dorsal interaction model, we posit that integrating semantic segmentation with depth estimation allows the system to plan and execute more adaptive and precise manipulations. To model the ventral and dorsal processing streams computationally, we utilize the Segment Anything Model 2 (SAM2) [15] for semantic segmentation, which corresponds to the ventral stream’s role in object identification. For estimating depth, we adopt the Depth Anything Model (DAM) [25], which mimics dorsal stream processing for spatial localization and motor planning. We expect that this dual-stream inspired integration will improve the robot’s ability to handle deformable, non-rigid objects in interaction with the human body, especially under conditions requiring fine motor adjustments and adaptive contact.
@@ IV-A Semantic extraction of garment and human foot
SAM2 and DAM are applied to estimate the state of socks and the human foot. SAM2 generates semantic mask images of socks and feet, and DAM generates depth images from a monocular RGB image. In general, segmentation models are vulnerable to external factors such as object deformation over time, occlusion, and varying lighting conditions [15]. During the sock dressing, various types of occlusions frequently occur, such as the sock itself covering the foot or the robot’s arms obscuring the target area. Furthermore, we utilize the SAM2 to mitigate the influence of variations in the color and appearance of the target object (e.g., socks or feet), as well as differences in the background surrounding the object. The images generated by SAM2 are black and white, with brightness values of 0 and 255, where the masked regions have a brightness value of 255. In parallel, the depth images of the images acquired from the robot’s camera are obtained using DAM[25]. The robot lacks perception in the depth direction when relying solely on two-dimensional images, so we employ DAM to infer the relative spatial relationship between the sock and the toes during insertion. Additionally, DAM facilitates estimating the vertical displacement of the arm required to successfully guide the sock over the heel region. Then, the mask depth images of the socks and feet are obtained by embedding the depth values of the depth image regions corresponding to the masked regions with a brightness value of 255. By adding depth information (z value) of the pixel value of the key points on the 2D coordinates of the image (x and y coordinates), 3D information on the image (x and y coordinates, and z value) is output.
@@
  • $D\in\mathbb{R}^{H\times W}$: The depth map obtained from the DAM.
@@ IV-B Visual and somatosensory attention mechanism
As a part of visual attention mechanism, Spatial Attention Recurrent Neural Network (SARNN) is applied for robustness to background change [27]. The image encoder of SARNN compresses and extracts features from vision, and the attention encoder extracts spatial attention points using the mask images as input. Multiple attention encoders enable the simultaneous estimation of attention points for multiple objects. To further enhance spatial understanding, we incorporate depth-aware attention. For each feature channel, a spatial softmax is applied to generate an attention map, which is used to weight the 2D spatial coordinates and depth map to calculate the expected 3D keypoints. This allows us to focus attention not only on 2D salient regions but also on the 3D geometric context, enabling robust keypoint localization even in occlusion and unknown situations.
@@ IV-C Hierarchical LSTM
The ${v}_{t}$ is split into modality-specific components:
@@ IV-D Training and inference phase
where $L_{angle}$, $L_{torque}$, $L_{tactile}$, $L_{img}$, and $L_{pt}$ represent the mean squared error (MSE) of the joint angles, the joint torques, the tactile values of the thumbs, the predicted image, and the attention point $pt=(x_{t},y_{t})$; $\alpha$, $\epsilon$$\beta$, $\gamma$, and $\delta$ represent the loss contributions set to 0.1, 0.1, 1.5, 1.0, and 0.2, respectively; Adam was used as the optimization model, and the model was trained for a total of 10,000 epochs.
@@ V-A Robot constitutions
In this study, we used a humanoid robot AIREC (Fig. 3) [14]. It integrates various sensing functions such as torque sensors, tactile sensors, and visual sensors, making it a platform for studying robot interaction with the real world. In addition, the control system is highly versatile, supporting various modes such as position, force, and impedance control, and parameters such as mass, damping, and spring can be customized. This adaptability allows the robot to perform tasks while improving safety, enhancing safety during contact with humans. Each arm of the robot has seven joints, and impedance controllers are used for the joints in the hand that may come into contact with the robot to soften them and improve safety. The robot incorporates hardware that facilitates reaching lower positions via waist bending.
@@ V-B Demonstration with PS controller
The robot was teleoperated using a PS controller. Hand positions were updated by adding xyz displacements from the controller inputs, and target joint angles were computed via inverse kinematics. Two motion patterns were used: parallel arm motion with identical displacements, and alternating arm motion in opposite directions. The parallel pattern, being intuitive and easy to reproduce, was primarily used for generating consistent demonstration data, while the alternating pattern was introduced to handle more diverse motions caused by friction or snagging.
@@ V-C Environment and dataset
In this study, a mannequin was used as a training subject. The specifications of the mannequin are 170 cm in height, 40 cm in shoulder width, and 19 cm in foot size. The exterior is made of urethane, and the interior is made of polystyrene foam. The socks were composed of rayon (66%), polyester (29%), silk (3%), and polyurethane (2%), with a size of 24–26 cm and a thickness of 0.9 mm. Stiffness was lower longitudinally (stress = 0.094 MPa, Young’s modulus = 0.103 MPa at 90.9% elongation) and higher transversely (stress = 0.497 MPa, Young’s modulus = 0.497 MPa at 100% elongation). As shown in Fig. 5, the mannequin was seated in a chair and foot was on a footrest. The seat height of the chair and the footrest were set to 45 cm. We assume a scenario in which a robot dresses socks for a person whose legs are out of bed and whose feet are off the floor. The participants maintained their natural sitting posture, and no quantitative assessment of posture was conducted.
The mannequin is lightweight, allowing its leg joints to bend easily. The mannequin’s feet are fixed with strings during training to apply the movements to real humans. Putting on socks requires balance,lower-body strength, and flexibility, which makes it a difficult movement for people with physical limitations. Collecting datasets in advance also places a burden on human participants. Therefore, a mannequin was used in the training phase for safety, and the adaptability of the model to humans was evaluated.
As shown in Fig. 4, since the feet position varies from person to person, we set the mannequin’s foot angle at three different positions—30°, 40°, and 50° from the horizontal plane. To balance the computation cost and training efficiency, we collected 4 samples for each, resulting in 12 samples for training. The algorithm was set to a frequency of 20 Hz, resulting in 210 time steps per dataset. The dataset includes the robot’s vision, the joint angles and torques of both arms’ joints 1-7, and the tactile information of the thumb, as shown in Fig. 3. The initial joint angles of the head were set to 0°, 37°, and 0° for head/joint1, head/joint2, and head/joint3, respectively. Similarly, the initial angles for the torso were set to 0°, -45°, and 100° for torso/joint1, torso/joint2, and torso/joint3, respectively. An impedance controller is applied to joint 7 to ensure safety in scenarios where contact with the human body may occur. The other joints employ position control, as actuation force is required to manipulate the sock effectively during the dressing process. In this study, the tactile information of the thumb is used to grasp a sock with the thumb and recognize the tightness of the sock.
@@ V-D2 Evaluation of generalization capability and robustness
To evaluate the generalization performance and robustness of the proposed model, we conducted experiments on 10 real humans and compared it with two existing models:
@@
The study involved ten participants with foot sizes between 23.0 cm and 26.5 cm. The model’s adaptability to environmental changes and individual differences in foot characteristics was assessed. Each method was evaluated in 50 trials per condition. Ethical approval was obtained from the Human Subject Research Ethics Review Committee of Waseda University (No. 2022-485).
@@ VI-A Ablation study on model components
In Table I, Dress-success and Insert-success denote the success rate of the complete dressing after toe-insertion and toe-insertion success rate, respectively. Removal of DAM or SKNet caused moderate performance degradation, while removal of Hierarchical LSTM or SAM2 and DAM caused a dramatic drop in dress-success rate, respectively. Without DAM, the model achieved a 85% success rate. However, one failure case was observed: after toe-insertion, when the arm was lowered along the surface of the feet, the toe got caught in the sock, making it difficult to continue movement.
Moreover, experiments validating the toe-insertion process revealed that the proposed model achieved the highest success rate compared with its ablated variants, as shown in Table I, indicating superior robustness in handling the insertion phase. However, four failures due to slight misalignments suggest that the system remains sensitive to positional differences. The insertion time was approximately 4 s. These results confirm the critical role of temporal modeling and the importance of spatial and depth-aware attention mechanisms.
@@
Model Variant Dress-success rate Insert-success rate
Ours 20/20 16/20
No DAM 17/20 6/20
No SKNet 16/20 8/20
No Hierarchical LSTM 1/20 2/20
No SAM and DAM 0/20 0/20
@@ VI-B Results of generalization capability and robustness
TABLE III: Maximum Tactile Forces for different foot sizes and phases
Foot Size [cm] toe [N] toe–heel [N] heel–ankle [N]
Ours ACT Ours ACT Ours ACT
24.0 19.08 10.88 20.94 12.83 21.83 12.59
25.0 19.20 9.11 20.87 12.94 20.93 12.88
26.0 19.19 9.01 20.96 13.63 22.16 13.63
Table II shows that our method achieved a success rate of 84% (42/50) in the trained background and 74% (37/50) in the untrained background. In contrast, ACT achieved 66% (33/50) and 0% (0/50), respectively, and DP failed to complete the dressing task (Fig. 6). Our method could mitigate environmental interference, such as lighting and shadow variations, although such factors still affect performance.
Fig. 7 shows the time series of thumb tactile values for different foot sizes. In Table III, the proposed model demonstrated a gradual increase in tactile force throughout the phases: $\approx 19$ N at the toe, $\approx 21$ N at toe–heel, and $\approx 22$ N at heel–ankle, exhibiting stable outputs across foot sizes. In contrast, the ACT model showed a different pattern, with minimal force at the toe phase and a maximum at the toe–heel. The ACT was more sensitive to size variations than the proposed method. The larger force at the heel–ankle phase in the proposed model was likely due to the deeper arm position and the wider hand spacing required to reach the heel area, enhancing tactile interaction. The consistent force across different foot sizes suggested that the proposed model effectively maintained an appropriate level of contact force to prevent excessive pressure during the dressing motion.
Human experiments revealed failure modes that were not observed in mannequin tests, primarily arising from morphological variability. Limited ankle flexibility and toenail snagging caused difficulties during toe insertion. Moreover, as shown in Table IV, foot size caused specific failures. The participants with larger feet experienced heel-catching failures due to premature arm lifting (Fig. 6). Consequently, although the system can accommodate a certain range of foot sizes, challenges remain in achieving robust performance across broader morphological variations. We would expand the method by leveraging a morpho-aware method [22].
A scene of attention points moving over the object area of semantic mask images is shown in Fig. 8. The attention points consistently track the object area. The proposed model was robust to unseen environments and individual anatomical variations, as shown in Fig.  9. The integration of semantic and somatosensory processing was effective, achieving efficient learning even with limited data. If the somatosensory modality were absent, the model would not function even with larger datasets. Fig. 10 shows the comparison results of the online performance between our model and ACT. This box plot shows the distribution of the number of successes out of 5 for 10 participants. We conducted a Wilcoxon signed rank test (Shapiro-Wilk test confirmed nonparametric) on the combined number of successes in the seen and unseen backgrounds. There was a statistically significant difference and large effect size between the proposed model and ACT (p-value << 0.01, Cohen’s dz = 1.11, 95% CI = [0.64, 1.93]). Although all models can perform well offline, ACT and DP trained on 12 datasets showed poor online performance due to inadequate force control, as shown in Table IV.
TABLE IV: Success rates corresponding to foot sizes.
@@
Figure 8: Attention points moving over the object area. Blue points and red points indicate current and future image attention key points, respectively.
@@
Stability and reproducibility are particularly important in the task of putting on socks. In DP, the consistency of the behavior was sometimes lost due to the probabilistic nature of behavior generation, making the movement of guiding the sock to the toe difficult and oscillatory. ACT tended to vary the arm trajectory. In an unknown background, ACT collided with the foot even in the first phase of reaching the foot. We assume that this is because the self-attention mechanism in Transformer has low sensitivity to local changes, which limits adaptation to differences due to foot size differences.
During the user study, socks sometimes caught on nails, making it difficult to continue the dressing motion. This necessitated motion replanning, highlighting the inherent complexity and unpredictability involved in sock-dressing assistance. Moreover, while our model performed better with limited training data, achieving a better trade-off between better generality and training cost remains future work.
@@
A novel multimodal imitation learning method, utilizing a hierarchical LSTM, for a humanoid robot assisting with close-fitting garment dressing was developed. Visual information is embedded with depth information in a semantic mask image, and the system recognizes the changing shapes of the foot and socks. The model trained on a mannequin’s foot demonstrated higher performance than the baseline in a subject experiment with 10 participants, taking into account various foot sizes, foot widths, and flexibility.